# On a class of interdiction problems with partition matroids: complexity and polynomial-time algorithms

### v3

@@ 3.2 Dynamic programming-based algorithm
21 begin
22       $\mathbf{r}^{*}_{k}\longleftarrow\left(r^{*}_{k-1,k^{\prime}}-\sum_{i\in F_{k-1}\cap L_{k^{\prime}}}x^{*}_{i}\right)_{k^{\prime}\in K_{l}}$
23       $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}\longleftarrow\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}^{*}_{k})$
@@
To implement dynamic programming, for each stage $k\in K_{f}$, we define a vector of states $\mathbf{r}_{k}\in\mathbb{Z}_{+}^{k_{l}}$ as the leader’s residual budgets at the beginning of stage $k$. Specifically, the sets of all possible states for stage $k=1$ and the remaining stages $k\in K_{f}\setminus\{1\}$ are defined, respectively, as:
@@
where we use $k^{\prime}\in K_{l}$ and $k\in K_{f}$ to denote the indices of the leader’s and the follower’s subsets. Also, for each stage $k\in K_{f}$, we introduce a set of actions
$A^{\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})=\Big{\{}\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}\in\{0,1\}^{|F_{k}|}:\;\sum_{i\,\in F_{k}\cap L_{k^{\prime}}}x_{i}\leq r_{k,k^{\prime}}\;\quad\forall k^{\prime}\in K_{l}\Big{\}},$ (15)
which corresponds to feasible leader’s decisions at stage $k$ for a given residual budget $\mathbf{r}_{k}\in R_{k}$. As a result, a new state, $\mathbf{r}_{k+1}$, can be defined as a function of the previous state, $\mathbf{r}_{k}$, and the previously selected blocking decision, $\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}$, i.e.,
$r_{k+1,k^{\prime}}=r_{k,k^{\prime}}-\sum_{i\in F_{k}\cap L_{k^{\prime}}}x_{i}\,\quad\forall k^{\prime}\in K_{l}.$ (16)
A pseudocode of the proposed DP-based algorithm is illustrated in Algorithm 1. Initially, at each stage $k\in\{k_{f},\ldots,1\}$ we compute a value function $V_{k}(\mathbf{r}_{k})$, which represents the minimum total loss the leader can achieve from a given state $\mathbf{r}_{k}\in R_{k}$ by following an optimal interdiction strategy; see lines 5 and 11 of Algorithm 1. In particular, in line 10, we express a new state, $\mathbf{r}_{k+1}$, as a function of $\mathbf{r}_{k}$ and $\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}$ by using equation (16). In lines 6 and 12, for each state $\mathbf{r}_{k}\in R_{k}$, we also identify an optimal leader’s decision, $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})$. It is noteworthy that, according to lines 6 and 12, Algorithm 1 identifies some optimal solution of [PMI] even when multiple optimal solutions exist. Finally, an optimal blocking decision, $\mathbf{x}^{*}$, is restored in lines 14–20 of Algorithm 1.
@@
We assert that for any $k\in K_{f}$, $k^{\prime}\in K_{l}$, and a pair of elements $i^{\prime}\in F_{k}\cap L_{k^{\prime}}$ and $i^{\prime\prime}\in F_{k}\cap L_{k^{\prime}}$ such that $i^{\prime}<i^{\prime\prime}$, there exists an optimal solution $\mathbf{x}^{*}\in\mathcal{I}_{X}$ satisfying $x^{*}_{i^{\prime}}\leq x^{*}_{i^{\prime\prime}}$. To this end, suppose that $x^{*}_{i^{\prime}}=1$ and $x^{*}_{i^{\prime\prime}}=0$. Then, it is always possible to switch the values of these two elements. Indeed, the newly obtained solution remains feasible by the definition of $\,\mathcal{I}_{X}$. Furthermore, it cannot increase the optimal leader’s objective function value, $\boldsymbol{\beta}^{\top}\mathbf{y}^{*}(\mathbf{x}^{*})$, as we block an element in $F_{k}$ with a potentially larger weight.
Using the outlined structural properties of $\mathbf{x}^{*}$, we conclude that the number of candidate optimal values for $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})$, $\mathbf{r}_{k}\in R_{k}$, can be reduced to
@@
Furthermore, since $r_{k,k^{\prime}}\in\{0,1,\ldots,l_{k^{\prime}}\}$ for any $k\in K_{f}$ and $k^{\prime}\in K_{l}$, the total number of possible combinations of $\mathbf{r}_{k}$ and $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})$ can be bounded from above as:
@@
Here, the first inequality holds because $\min\{r_{k,k^{\prime}};|F_{k}\cap L_{k^{\prime}}|\}\leq r_{k,k^{\prime}}$. The last inequality follows from the inequality of arithmetic and geometric means and our assumption that $\sum_{k^{\prime}\in K_{l}}l_{k^{\prime}}\leq n$. We conclude that there are at most $O(n^{2k_{l}})$ possible combinations of $\mathbf{r}_{k}$ and $\mathbf{x}^{*\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})$ for $k\in K_{f}$.
@@
$V_{k}(\mathbf{r}_{k})=\min_{\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}\in A^{\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})}\Big{\{}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}})\Big{\}},$ (18)
where $k=k_{f}$ and $\mathbf{r}_{k_{f}}\in R_{k_{f}}$. There are $O(n^{2k_{l}})$ possible combinations of $\mathbf{r}_{k_{f}}$ and $\mathbf{x}^{*\mbox{\tiny(k}_{f}\mbox{\tiny)}}(\mathbf{r}_{k_{f}})$ and, also, for any fixed optimal decision $\mathbf{x}^{*\mbox{\tiny(k}_{f}\mbox{\tiny)}}(\mathbf{r}_{k_{f}})$, the objective function in (18) can be computed in $O(|F_{k_{f}}|\log|F_{k_{f}}|)$ operations by the greedy algorithm. Therefore, the total number of operations required to compute $V_{k_{f}}(\mathbf{r}_{k_{f}})$ is of the order
@@
$V_{k}(\mathbf{r}_{k})=\min_{\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}}\in A^{\mbox{\tiny(k}\mbox{\tiny)}}(\mathbf{r}_{k})}\Big{\{}\psi_{k}(\mathbf{x}^{\mbox{\tiny(k}\mbox{\tiny)}})+V_{k+1}(\mathbf{r}_{k+1})\Big{\}};$
recall line 11. Analogously to the value function at stage $k_{f}$, computing $V_{k}(\mathbf{r}_{k})$ requires $O(n^{2k_{l}}|F_{k}|)$ operations. Finally, from lines 1414141420202020 of Algorithm 1 we observe that an optimal solution $\mathbf{x}^{*}$ can be restored in $O(n)$ operations.
@@ 3.3 Greedy algorithm
and the feasible sets $\mathcal{I}_{X}$ and $\mathcal{I}_{Y}(\mathbf{x})$, $\mathbf{x}\in\mathcal{I}_{X}$, are given by equations (2a) and (2b), respectively. Given that $\mathbf{x}$ is binary, we demonstrate that the leader’s objective function, $\psi(\mathbf{x})$, can be viewed as a monotone and submodular set function. The following result holds.
@@
    $\psi(\mathbf{x})$ is monotone, i.e., for any $\mathbf{x}^{\prime}\in\{0,1\}^{n}$ and $\mathbf{x}^{\prime\prime}\in\{0,1\}^{n}$, if $\mathbf{x}^{\prime}\leq\mathbf{x}^{\prime\prime}$, then
    $\psi(\mathbf{x}^{\prime})\leq\psi(\mathbf{x}^{\prime\prime}).$
@@ 8       $\mathbf{x}\longleftarrow\mathbf{x}+\mathbf{e}_{i^{*}}$
Given that the leader’s objective function, $\psi(\mathbf{x})$, is non-increasing, a myopic approach to approximately solve [PMI] is to consider a simple greedy algorithm for the leader; we refer to Algorithm 2 for its pseudocode. The algorithm starts with an empty independent set by setting $\mathbf{x}=\mathbf{0}$ (line 3). In lines 5–10, we increment the current blocking decision $\mathbf{x}$, while preserving feasibility and maximizing the decrease in the leader’s objective function value. In the following, whenever the argmin operator in line 7 yields multiple optimal solutions, we break ties by selecting an index $i^{*}$ with the largest possible weight. Finally, Algorithm 2 terminates once the current independent set can no longer be expanded.
@@
We demonstrate that, under Assumption A1, Algorithm 2 yields an exact solution for [PMI] when the follower’s feasible set is defined by a uniform matroid, i.e., $k_{f}=1$. This result is somewhat surprising, as [PMI] with $k_{f}=1$ constitutes a submodular minimization problem, which typically cannot be solved using a greedy algorithm.
@@
Next, using equation (20), we observe that for any $\mathbf{x}\in\mathcal{I}_{X}$ and $i\in N(\mathbf{x})$
@@
recall line 7 of Algorithm 2. Furthermore, Assumption A1 implies that for all $i\geq j(\mathbf{x},f)$
$\beta_{j(\mathbf{x},f+1)}-\beta_{i}<0.$
Let us define
@@
Then, in the first stage, Algorithm 2 subsequently selects elements $i\in N_{+}(\mathbf{x})$ with the largest possible weight $\beta_{i}$; recall (21). In the second stage, whenever $N_{+}(\mathbf{x})=\emptyset$, the algorithm continues by expanding $\mathbf{x}$ with elements $i\in N_{-}(\mathbf{x})$, also maximizing $\beta_{i}$; recall that, in case of multiple solutions in line 7, we break ties by selecting an index with the largest possible weight. We conclude that Algorithm 2 for $k_{f}=1$ implicitly solves the maximum weight independent set problem for the leader given by:
$\max_{\mathbf{x}\in\,\mathcal{I}_{X}}\sum_{i=1}^{n}\beta_{i}x_{i}.$ (22)
Finally, we show that the optimal solution of (22) is also optimal for [PMI] with $k_{f}=1$. First, under Assumption A1, the optimal solution of (22) is unique. Furthermore, according to Theorem 2, the dual reformulation of [PMI] with $k_{f}=1$ reads as:
$\begin{gathered}\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\,\min_{j\in\{0,\ldots,n\}}\Big{\{}\beta_{j}f\;+\;\sum_{i=j+1}^{n}(\beta_{i}-\beta_{j})(1-x_{i})\Big{\}}=\\ \min_{j\in\{0,\ldots,n\}}\Big{\{}\beta_{j}f\;+\;\sum_{i=j+1}^{n}(\beta_{i}-\beta_{j})-\max_{\mathbf{x}\in\mathcal{I}_{X}}\sum_{i=j+1}^{n}(\beta_{i}-\beta_{j})x_{i}\Big{\}}.\end{gathered}$
By design of the greedy algorithm for matroids, the optimal solution of (22) is also optimal for each subproblem
$\max_{\mathbf{x}\in\mathcal{I}_{X}}\sum_{i=j+1}^{n}(\beta_{i}-\beta_{j})x_{i},$
where $j\in\{0,\ldots,n\}$. Hence, the optimal solution of (22) is also optimal for [PMI] with $k_{f}=1$ and the result follows. ∎

### v4

@@ 20 $\mathbf{x}^{\text{\tiny(1)}}\longleftarrow\mathbf{x}^{\text{\tiny(1)}}(\mathbf{r}^{*}_{1})$
[…]
23       $\mathbf{x}^{*\text{\tiny({k})}}\longleftarrow\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}^{*}_{k})$
@@
In the following, $k^{\prime}\in K_{l}$ and $k\in K_{f}$ are used to denote the indices of the leader’s and the follower’s subsets. To implement dynamic programming, for each stage $k\in K_{f}$, we define a vector of states $\mathbf{r}_{k}\in\mathbb{Z}_{+}^{k_{l}}$ as the leader’s residual budgets at the beginning of stage $k$. Specifically, the sets of all possible states for stage $k=1$ and the remaining stages $k\in K_{f}\setminus\{1\}$ are defined, respectively, as:
@@
where $r_{k,k^{\prime}}$ represents the residual budget of the leader related to the constraint $\sum_{i\in L_{k^{\prime}}}x_{i}\leq l_{k^{\prime}}$. Also, for each stage $k\in K_{f}$, we introduce a set of actions
$A^{\text{\tiny({k})}}(\mathbf{r}_{k})=\Big{\{}\mathbf{x}^{\text{\tiny({k})}}\in\{0,1\}^{|F_{k}|}:\;\sum_{i\,\in F_{k}\cap L_{k^{\prime}}}x_{i}\leq r_{k,k^{\prime}}\;\quad\forall k^{\prime}\in K_{l}\Big{\}},$ (15)
which corresponds to feasible leader’s decisions at stage $k$ for a given residual budget $\mathbf{r}_{k}\in R_{k}$. As a result, a new state, $\mathbf{r}_{k+1}$, can be defined as a function of the previous state, $\mathbf{r}_{k}$, and the previously selected blocking decision, $\mathbf{x}^{\text{\tiny({k})}}$, i.e.,
$\mathbf{r}_{k+1}=\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}}):=\Big{(}r_{k,k^{\prime}}-\sum_{i\in F_{k}\cap L_{k^{\prime}}}x_{i}\Big{)}_{k^{\prime}\in K_{l}}.$ (16)
A pseudocode of the proposed DP-based algorithm is illustrated in Algorithm 1. Initially, at each stage $k\in\{k_{f},\ldots,1\}$, we compute a value function $V_{k}(\mathbf{r}_{k})$, which represents the minimum total loss the leader can achieve from a given state $\mathbf{r}_{k}\in R_{k}$ by following an optimal interdiction strategy; see lines 5 and 11 of Algorithm 1. In particular, in line 10, we express a new state, $\mathbf{r}_{k+1}$, as a function of $\mathbf{r}_{k}$ and $\mathbf{x}^{\text{\tiny({k})}}$ by using equation (16). In lines 6 and 12, for each state $\mathbf{r}_{k}\in R_{k}$, we also identify an optimal leader’s decision, $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})$. It is noteworthy that, according to lines 6 and 12, Algorithm 1 identifies some optimal solution of [PMI] even when multiple optimal solutions exist. Finally, an optimal blocking decision, $\mathbf{x}^{*}$, is restored in lines 15–20 of Algorithm 1.
@@
We assert that, for any $k\in K_{f}$, $k^{\prime}\in K_{l}$, and a pair of elements $i^{\prime}\in F_{k}\cap L_{k^{\prime}}$ and $i^{\prime\prime}\in F_{k}\cap L_{k^{\prime}}$ such that $i^{\prime}<i^{\prime\prime}$, there exists an optimal solution $\mathbf{x}^{*}\in\mathcal{I}_{X}$ satisfying $x^{*}_{i^{\prime}}\leq x^{*}_{i^{\prime\prime}}$. To this end, suppose that $x^{*}_{i^{\prime}}=1$ and $x^{*}_{i^{\prime\prime}}=0$. Then, it is always possible to switch the values of these two elements. Indeed, the newly obtained solution remains feasible by the definition of $\,\mathcal{I}_{X}$. Furthermore, it cannot increase the optimal leader’s objective function value, $\boldsymbol{\beta}^{\top}\mathbf{y}^{*}(\mathbf{x}^{*})$, as we block an element in $F_{k}$ with a potentially larger weight.
Using the outlined structural properties of $\mathbf{x}^{*}$, we conclude that the number of candidate optimal values for $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})$, $\mathbf{r}_{k}\in R_{k}$, can be reduced to
@@
Furthermore, since $r_{k,k^{\prime}}\in\{0,1,\ldots,l_{k^{\prime}}\}$ for any $k\in K_{f}$ and $k^{\prime}\in K_{l}$, the total number of possible combinations of $\mathbf{r}_{k}$ and $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})$ can be bounded from above as:
@@
Here, the first inequality holds because $\min\{r_{k,k^{\prime}};|F_{k}\cap L_{k^{\prime}}|\}\leq r_{k,k^{\prime}}$. The last inequality follows from the inequality of arithmetic and geometric means and our assumption that $\sum_{k^{\prime}\in K_{l}}l_{k^{\prime}}\leq n$. We conclude that there are at most $O(n^{2k_{l}})$ possible combinations of $\mathbf{r}_{k}$ and $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})$ for $k\in K_{f}$.
@@
$V_{k}(\mathbf{r}_{k})=\min_{\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})}\Big{\{}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})\Big{\}},$ (18)
where $k=k_{f}$ and $\mathbf{r}_{k_{f}}\in R_{k_{f}}$. There are $O(n^{2k_{l}})$ possible combinations of $\mathbf{r}_{k_{f}}$ and $\mathbf{x}^{*\text{\tiny({k}}_{f}\text{\tiny)}}(\mathbf{r}_{k_{f}})$ and, also, for any fixed optimal decision $\mathbf{x}^{*\text{\tiny({k}}_{f}\text{\tiny)}}(\mathbf{r}_{k_{f}})$, the objective function in (18) can be computed in $O(|F_{k_{f}}|\log|F_{k_{f}}|)$ operations by the greedy algorithm. Therefore, the total number of operations required to compute $V_{k_{f}}(\mathbf{r}_{k_{f}})$ is of the order
@@
$V_{k}(\mathbf{r}_{k})=\min_{\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})}\Big{\{}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})+V_{k+1}\big{(}\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\big{)}\Big{\}};$
recall line 11. Analogously to the value function at stage $k_{f}$, computing $V_{k}(\mathbf{r}_{k})$ requires $O(n^{2k_{l}}|F_{k}|)$ operations, since the value function
$V_{k+1}(\mathbf{r}_{k+1})=V_{k+1}\big{(}\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\big{)}$
for each $\mathbf{r}_{k+1}\in R_{k+1}$ is precomputed at stage $k+1$. Finally, from lines 1414141420202020 of Algorithm 1 we observe that an optimal solution $\mathbf{x}^{*}$ can be restored in $O(n)$ operations.
@@ 3.3 Greedy algorithm
and the feasible sets $\mathcal{I}_{X}$ and $\mathcal{I}_{Y}(\mathbf{x})$, $\mathbf{x}\in\mathcal{I}_{X}$, are given by equations (2a) and (2b), respectively. Given that $\mathbf{x}$ is binary, we demonstrate that the leader’s objective function, $\psi(\mathbf{x})$, can be viewed as a monotone non-increasing and submodular set function. The following result holds.
@@
    $\psi(\mathbf{x})$ is monotone non-increasing, i.e., for any $\mathbf{x}^{\prime}\in\{0,1\}^{n}$ and $\mathbf{x}^{\prime\prime}\in\{0,1\}^{n}$, if $\mathbf{x}^{\prime}\leq\mathbf{x}^{\prime\prime}$, then
    $\psi(\mathbf{x}^{\prime})\geq\psi(\mathbf{x}^{\prime\prime}).$
@@ 7       $\mathbf{x}\longleftarrow\mathbf{x}+\mathbf{e}_{i^{*}}$
As outlined in Section 1.3, given that the leader’s objective function, $\psi(\mathbf{x})$, is non-increasing, it is natural to consider a simple greedy algorithm for the leader in the context of [PMI]; we refer to Algorithm 2 for its pseudocode. The algorithm starts with an empty independent set by setting $\mathbf{x}=\mathbf{0}$ (line 3). In lines 5–9, we increment the current blocking decision $\mathbf{x}$, while preserving feasibility and maximizing the decrease in the leader’s objective function value (ties broken arbitrarily). Finally, Algorithm 2 terminates once the current independent set can no longer be expanded.
@@
Next, we demonstrate that, under Assumption A1, Algorithm 2 yields an exact solution for [PMI] when the follower’s feasible set is defined by a uniform matroid, i.e., $k_{f}=1$. This result is somewhat surprising, as [PMI] with $k_{f}=1$ constitutes a submodular minimization problem, which typically cannot be solved using a greedy algorithm.
@@
Next, using equation (20), we observe that, for any $\mathbf{x}\in\mathcal{I}_{X}$ and $i\in N(\mathbf{x})$,
@@
recall line 6 of Algorithm 2. Let us define
@@
Then, in the first stage, Algorithm 2 subsequently selects elements $i\in N_{+}(\mathbf{x})$ with the largest possible weight $\beta_{i}$; recall (21) and Assumption A1. In the second stage, whenever $N_{+}(\mathbf{x})=\emptyset$, the algorithm continues by expanding $\mathbf{x}$ with elements $i\in N_{-}(\mathbf{x})$, breaking ties arbitrarily.
It suffices to show that a solution $\mathbf{x}_{g}\in\mathcal{I}_{X}$ obtained by Algorithm 2 solves [PMI] with $k_{f}=1$. In this regard, we note that, by construction, $\mathbf{x}_{g}$ solves the following maximum weight independent set problem:
$\max_{\mathbf{x}\in\,\mathcal{I}_{X}}\sum_{i=j(\mathbf{x}_{g},f)}^{n}\beta_{i}x_{i}$
and, consequently, by leveraging (20), we have:
$\psi(\mathbf{x}_{g})=\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\sum_{i=j(\mathbf{x}_{g},f)}^{n}\beta_{i}(1-x_{i}).$ (22)
Assume, for the sake of contradiction, that
$\psi(\mathbf{x}_{g})>\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\psi(\mathbf{x}).$ (23)
According to Theorem 2, we have:
$\begin{gathered}\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\psi(\mathbf{x})=\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\,\min_{j\in\{0,\ldots,n\}}\Big{\{}\beta_{j}f\;+\;\sum_{i=j+1}^{n}(\beta_{i}-\beta_{j})(1-x_{i})\Big{\}}=\\ \min_{j\in\{0,\ldots,n\}}\Big{\{}\beta_{j}f\;+\;\min_{\mathbf{x}\in\mathcal{I}_{X}}\sum_{i=j+1}^{n}(\beta_{i}-\beta_{j})(1-x_{i})\Big{\}}.\end{gathered}$
Let $\mathbf{x}^{*}\in\mathcal{I}_{X}$ be a solution corresponding to the maximum weight independent set in $X$; this solution is unique by Assumption A1. Clearly,
$\psi(\mathbf{x}_{g})=\psi(\mathbf{x}^{*}),$
since the choice of elements with indices $i<j(\mathbf{x}_{g},f)$ does not affect the leader’s objective function value in (22). Furthermore, by design of the greedy algorithm for matroids, $\mathbf{x}^{*}$ is also optimal for each subproblem
$\min_{\mathbf{x}\in\mathcal{I}_{X}}\sum_{i=j+1}^{n}(\beta_{i}-\beta_{j})(1-x_{i}),$
where $j\in\{0,\ldots,n\}$. Hence,
$\psi(\mathbf{x}_{g})=\psi(\mathbf{x}^{*})=\min_{\mathbf{x}\in\,\mathcal{I}_{X}}\psi(\mathbf{x}),$
that contradicts inequality (23). This observation concludes the proof. ∎